# Attention U-Net for all-sky continuous gravitational wave searches

### v1

@@ I Introduction
The amplitude of CWs is expected to be significantly weaker than the noise level of current gravitational wave detectors. To achieve sufficient sensitivity, integrating a long duration of data is required to increase the signal-to-noise ratio (SNR). However, the most sensitive method, template-based coherent matched filtering, demands an increasing number of templates for longer observation times, with computational cost scaling as $\sim T^{n}$, where $n\sim 5$ to 6 or higher. This limits coherent searches to a few weeks of data. Alternatively, semi-coherent methods divide the total observation time into shorter segments, coherently analyzing each segment and combining the segment detection statistics incoherently. Despite their reduced computational burden, semi-coherent methods remain challenging and computationally bound for all-sky CW searches because of the vast parameter space involved. To overcome these computational limitations, machine learning offers a promising approach for effectively navigating the extensive parameter space.
@@
For all-sky CW searches, all models address signal detection as a classification problem. (Dreissigacker et al., 2019; Dreissigacker and Prix, 2020) trained a ResNet on 11.6 days of data with Gaussian noise. (Yamamoto et al., 2022) used a CNN to detect signals with Gaussian and sinusoidal line noise. (Joshi and Prix, 2023, 2024) developed a CNN with residual blocks to look for the presence of signal on 10 days of data. (Joshi and Prix, 2025) trained a transformer-based model on 1 day of data. These represent the only end-to-end machine learning-based all-sky search models to date. Among them, Joshi and Prix (2024) achieved the highest sensitivity across the largest parameter space explored so far.
In this work, we employ an attention U-Net (Oktay et al., 2018), a CNN commonly used for image segmentation, to develop an end-to-end all-sky CW search model. We treat the input data as images and train the U-Net to map noisy inputs to denoised signal outputs, using the total power of the output image as the detection statistic. Five models were trained for all-sky CW signals at 20, 200, 500, 1000, and the entire 20–1000 Hz band, respectively. Our model trained at 20 Hz achieves the best sensitivity, with a sensitivity depth (see Sect. II) of $\mathcal{D}^{90\%}=29.97\pm 0.24\,\mathrm{Hz}^{-1/2}$ at 90% detection efficiency, and $\mathcal{D}^{90\%}=18.63\pm 0.24\,\mathrm{Hz}^{-1/2}$ for the model trained on the entire 20–1000 Hz band. The sensitivities are comparable to the results reported by (Joshi and Prix, 2024), using less than 50% of the training time and data, requiring approximately 5 days for training. This serves as a proof of concept that the attention U-Net is a suitable and effective approach for CW searches. For the 20 Hz model, we extended the observation time, achieving a sensitivity depth of $\mathcal{D}^{90\%}=35.86\pm 0.38\,\mathrm{Hz}^{-1/2}$ with $T\approx 21$ days, with sensitivity scaling as $T^{0.28\pm 0.01}$, comparable to the $T^{0.25}$ scaling of semi-coherent searches (Wette, 2012). The training time scales linearly with observation time, more efficient when compare the to semi-coherent method $\sim T^{2}$ dependece (Prix and Shaltev, 2012). When tested on datasets with time gaps, the model, trained on continuous data, exhibits robustness under real data conditions, with sensitivity dependence on duty factor scaling as $\eta^{0.53\pm 0.02}$. Additionally, we analyzed the sensitivity dependence of the trained models on sky location to investigate their limitations.
The remainder of this article is organized as follows. Section II describes how we simulate the CW signals. Section III details the attention U-Net architecture and the neural network training procedure. Section IV presents the performance results of the trained attention U-Net on test data. Section V discusses potential improvements and future research directions.
@@ II Waveform Models
where $n(t)$ represents the detector noise and $h(t)$ denotes the CW signal. The CW signal from an isolated, rapidly spinning, non-axisymmetric neutron star (NS) is given by
@@
For CW data analysis, we convert the time series into the Short Fourier Transform (SFT) format by dividing the data into segments. Under the assumption of a slowly varying CW signal, SFT data enable signal detection while significantly reducing dataset size. Although detector noise in practice exhibits non-Gaussian properties, we assume Gaussian noise for simplicity. We use PyFstat (Keitel et al., 2021) to generate pure CW signals in SFT format, segmenting the data into time spans of $T_{\text{SFT}}=144{,}000$ s (4 hours) per segment. Gaussian noise is generated directly in the Fourier domain as
@@
We use $\sqrt{S_{n}}/h_{0}$ to represent the noise level of the injected signal in our training and test data, as the ability to detect the signal depends on its amplitude relative to the detector’s noise amplitude. In this work, we consider only the first-order spindown term in Eq. (3). The parameter space defining the signals is listed in Table 1. We assess the neural network’s performance for an all-sky search with an observation time of T=921,600T=921{,}600 s (approximately 10.67 days), chosen to approximate 10 days and to yield $2^{n}$ time segments to optimize the neural network’s architecture.
@@ III Methods
Similar to (Dreissigacker and Prix, 2020; Joshi and Prix, 2024), we use the real and imaginary parts of SFT data as a two-dimensional input. Combining data from the H1 and L1 detectors, we obtain four channels: real and imaginary components for each detector. We perform a Fourier transform every 14,400 s on data spanning a total observation time of T=921,600T=921{,}600 s (approximately 10.67 days), yielding 64 time segments along the time axis. The frequency axis is cropped to retain 512 frequency bins, corresponding to a bandwidth of 35.56 mHz, sufficient to accommodate Doppler modulations due to Earth’s motion over a 10-day observation period. To enhance robustness, we then randomly shift the frequency axis, ensuring the CW signal remains within the selected range without always appearing at the center. The resulting preprocessed data have a shape of $(512,64,4)$, where the dimensions are chosen to approximate 10 days and to yield powers of 2 to optimize the U-Net’s architecture. Finally, we normalize the data to the range [1,1][-1,1].
@@ III.4 Detection Statistic and Detectability
In addition to the loss function, we monitored the training process with $p_{\mathrm{det}}$, which measures the fraction of noisy CW signal samples correctly identified by the detection statistic at a fixed false alarm rate ($p_{\mathrm{fa}}$). Using Eq. 3 from Joshi and Prix (2024), we set $p_{\mathrm{fa}}$ to 1% per 50 mHz. For our data, spanning a 35.56 mHz frequency band per image, we calculated $p_{\mathrm{fa}}=0.71\%$. We selecte the model with the highest $p_{\mathrm{det}}$ on the validation set as the final model, since $p_{\mathrm{det}}$ is the metric for evaluating the model’s ability to identify CW signals in the presence of noise.
@@ IV Results
Figure 1: Loss (top) and detection probability $p_{\mathrm{det}}$ (at a fixed $p_{\mathrm{fa}}=1\%$ per $50\,\mathrm{mHz}$) (bottom) during training for the 20 Hz model on the training dataset (black) and validation dataset (colored) with different noise level $\sqrt{S_{n}}/h_{0}$.
@@ IV.1 Denoising
Figure 2: Denoising performance of the trained attention U-Net at 20 Hz (left) and 500 Hz (right) on H1 spectrogram spanning 921,600 s. The attention U-Net takes a noisy input with noise level $\sqrt{S_{n}}/h_{0}=20\,\mathrm{Hz}^{-1}$ (top) and produces a denoised output (middle), which closely resembles the ground truth signal (bottom), recovering the major signal features.
@@
Figure 3: Denoising performance of the trained attention U-Net at 20 Hz (left) and 500 Hz (right) on H1 spectrogram spanning 921,600 s with time gaps (duty factor = 0.75). Despite being trained on data without time gaps, the attention U-Net successfully denoises the noisy input with noise level $\sqrt{S_{n}}/h_{0}=20\,\mathrm{Hz}^{-1}$ (top) to produce an output (middle) that closely resembles the ground truth signal (bottom), recovering the major signal features.
@@ IV.2 Sensitivity
We evaluated $p_{\mathrm{det}}$ across various noise levels $\sqrt{S_{n}}/h_{0}$ using 2,000 test samples (1,000 noisy signals and 1,000 pure noise) to determine the noise level at which a 90% detection probability is achieved ($\mathcal{D}^{90\%}$). Figure 4 shows the $p_{\mathrm{det}}$ of the 20 Hz model at different noise levels (blue curve). We also included the $p_{\mathrm{det}}$ of a matched-filter search using Weave from (Joshi and Prix, 2024), scaled to match the timespans in this study, as a benchmark (black curve). Our model achieves a sensitivity depth of $\mathcal{D}^{90\%}=29.97\pm 0.24\,\mathrm{Hz}^{-1/2}$. To compare with (Joshi and Prix, 2024), we scaled their $\mathcal{D}^{90\%}$ values using the relation $\mathcal{D}^{90\%}\propto T^{0.25}$, adjusting from their 10-day observation period to our 921,600 s ($\approx$ 10.67 days). This yields the sensitivities of $29.78\,\mathrm{Hz}^{-1/2}$ at 20 Hz. The sensitivity of our attention U-Net is comparable to the scaled sensitivity from (Joshi and Prix, 2024) at 20 Hz.
Figure 4: Detection probability $p_{\mathrm{det}}$ (at a fixed $p_{\mathrm{fa}}=1\%$ per $50\,\mathrm{mHz}$) versus noise level $\sqrt{S_{n}}/h_{0}$ for the 20 Hz model. The black curve shows the performance of the Weave matched-filter method, rescaled from (Joshi and Prix, 2024). The blue curve represents the attention U-Net’s performance, with a 90% sensitivity depth $\mathcal{D}^{90\%}=29.97\pm 0.24\,\mathrm{Hz}^{-1/2}$.
The performance of all models is summarized in Table 2, which reports the sensitivity depth $\mathcal{D}^{90\%}$ of our trained model, the rescaled sensitivity depth using the Weave matched-filter method $\mathcal{D}_{\mathrm{MF}}^{90\%}$ from (Joshi and Prix, 2024), and the detectability of our trained model at a noise level equals to $\sqrt{S_{n}}/h_{0}=\mathcal{D}_{\mathrm{MF}}^{90\%}$. Our attention U-Net achieves comparable sensitivity at 20 Hz and 200 Hz but lower sensitivity at 500 Hz, 1000 Hz, and acroos the 20–1000 Hz range compared to (Joshi and Prix, 2024).
@@
Table 2: Sensitivity depth at 90% detection probability ($\mathcal{D}^{90\%}$) for the trained attention U-Net, compared to the matched-filter sensitivity depth ($\mathcal{D}^{90\%}_{\mathrm{MF}}$) from (Joshi and Prix, 2024), and the detection probability of the trained attention U-Net at noise level $\sqrt{S_{n}}/h_{0}=\mathcal{D}_{\mathrm{MF}}^{90\%}$.
@@ IV.3 Generalization on Duty Factor and Total Observation Time
To study the dependence of sensitivity on the duty factor $\eta$, we evaluated the performance of the 20 Hz model on test data with time gaps, simulating realistic observation conditions. Figure 5 shows the variation of $\mathcal{D}^{90\%}$ with duty factors ranging from 0.5 to 1.0, where typical observation runs have duty factors of approximately 0.6–0.7. Fitting $\mathcal{D}^{90\%}$ to a power law reveals a scaling of $\mathcal{D}^{90\%}\propto\eta^{0.53\pm 0.02}$, with $\mathcal{D}^{90\%}=22.97\pm 0.26$ at $\eta=0.6$.
Figure 5: Sensitivity depth $\mathcal{D}^{90\%}$ versus duty factor $\eta$ for the 20 Hz model. The sensitivity scales as $\mathcal{D}^{90\%}\propto\eta^{0.53\pm 0.02}$.
Our neural network was initially trained on data with $T\approx$ 10.67 days. To maximize sensitivity for CW signal detection, we aim to use as much data as possible. To investigate the impact of $T$ on the sensitivity, we trained the 20 Hz model with observation time of 460,800 s, 921,600 s, 1,382,400 s, and 1,832,000 s (approximately 5, 10, 16, and 21 days, respectively). Figure 6 shows the resulting $\mathcal{D}^{90\%}$ values, achieving $\mathcal{D}^{90\%}=35.86\pm 0.38\,\mathrm{Hz}^{-1/2}$ at $T\approx$ 21 days. The sensitivity depth scales as $\mathcal{D}^{90\%}\propto T^{0.28\pm 0.01}$, similar to the $T^{0.25}$ scaling observed in semicoherent search methods. Notably, the training time for our attention U-Net scales linearly with $T$, as the number of frequency bins is fixed. For sufficiently long $T$, such as a year of data, we have to increase the number of time bins by a factor of 35 and frequency bins by a factor of 7 to cover the entire signal for the parameter range defined in Table. 1. As a result, the overall scaling remains lower than that of semicoherent searches, which scale as $\sim T^{2}$ for a fixed coherence time (Prix and Shaltev, 2012). This demonstrates that the attention U-Net architecture is feasible to scale up to longer observation time for better performance.
Figure 6: Sensitivity depth $\mathcal{D}^{90\%}$ versus total observation time $T$ for the 20 Hz model, evaluated at $T$ = 460,800 s, 921,600 s, 1,382,400 s, 1,832,000 s (approximately 5, 10, 16, and 21 days). The model achieves $\mathcal{D}^{90\%}=35.86\pm 0.38\,\mathrm{Hz}^{-1/2}$ at $T\approx$ 21 days. The fitted power law shows $\mathcal{D}^{90\%}\propto T^{0.28\pm 0.01}$.

### v2

@@ I Introduction
The amplitude of CWs is expected to be significantly weaker than the noise amplitude of current gravitational wave detectors. To achieve sufficient sensitivity, integrating a long duration of data is required to increase the signal-to-noise ratio (SNR). However, the most sensitive method, template-based coherent matched filtering, demands an increasing number of templates for longer observation times, with computational cost scaling as $\sim T^{n}$, where $n\sim 5$ to 6 or higher. This limits coherent searches to a few weeks of data. Alternatively, semi-coherent methods divide the total observation time into shorter segments, coherently analyzing each segment and combining the segment detection statistics incoherently. Despite their reduced computational burden, semi-coherent methods remain challenging and computationally bound for all-sky CW searches because of the vast parameter space involved. To overcome these computational limitations, machine learning offers a promising approach for effectively navigating the extensive parameter space.
@@
For all-sky CW searches, all models address signal detection as a classification problem. (Dreissigacker et al., 2019; Dreissigacker and Prix, 2020) trained a ResNet on $\approx 11.57$ days of data with Gaussian noise. (Yamamoto et al., 2022) used a CNN to detect signals with Gaussian and sinusoidal line noise. (Joshi and Prix, 2023, 2024) developed a CNN with residual blocks to look for the presence of signal on 10 days of data. (Joshi and Prix, 2025) trained a transformer-based model on 1 day of data. These represent the only end-to-end machine learning-based all-sky search models to date. Among them, (Joshi and Prix, 2024) achieved the highest sensitivity across the largest parameter space explored so far.
In this work, we employ an attention U-Net (Oktay et al., 2018), a CNN commonly used for image segmentation, to develop an end-to-end all-sky CW search model. We treat the input data as images and train the U-Net to map noisy inputs to denoised signal outputs, using the mean power of the output image as the detection statistic. Five models were trained for all-sky CW signals at 20, 200, 500, 1000, and the entire 20–1000 Hz band, respectively. Our model trained at 20 Hz achieves the best sensitivity, with a sensitivity depth (see Sect. II) of $\mathcal{D}^{90\%}=29.97\pm 0.24\,\mathrm{Hz}^{-1/2}$ at 90% detection efficiency, and $\mathcal{D}^{90\%}=18.63\pm 0.24\,\mathrm{Hz}^{-1/2}$ for the model trained on the entire 20–1000 Hz band. The sensitivities are comparable to the results reported by (Joshi and Prix, 2024), using less than 50% of the training time and data, requiring approximately 5 days for training. This serves as a proof of concept that the attention U-Net is a suitable and effective approach for CW searches. For the 20 Hz model, we extended the observation time, achieving a sensitivity depth of $\mathcal{D}^{90\%}=35.86\pm 0.38\,\mathrm{Hz}^{-1/2}$ with $T\approx 21$ days, with sensitivity scaling as $T^{0.28\pm 0.01}$, comparable to the $T^{0.25}$ scaling of semi-coherent searches (Wette, 2012). The training time scales linearly with observation time, more efficient when compare the to semi-coherent method $\sim T^{2}$ dependece (Prix and Shaltev, 2012). When tested on datasets with time gaps, the model, trained on continuous data, exhibits robustness under real data conditions, with sensitivity dependence on duty factor scaling as $\eta^{0.53\pm 0.02}$. Additionally, we analyzed the sensitivity dependence of the trained models on sky location to investigate their limitations.
This article is organized as follows. Section II outlines the CW waveform models. Section III describes the attention U-Net architecture, training process, and detection statistic for assessing detectability. Section IV evaluates the performance of the trained attention U-Net models on test data. Finally, Section V explores potential improvements and future research directions.
@@ II Waveform Models
where $n(t)$ represents the detector noise and $h(t)$ denotes the CW signal. The CW signal from an isolated, rapidly spinning, non-axisymmetric NS is given by
@@
For CW data analysis, we convert the time series into the Short Fourier Transform (SFT) format by dividing the data into segments. Under the assumption of a slowly varying CW signal, SFT data enable signal detection while significantly reducing the size of the dataset. We use PyFstat (Keitel et al., 2021) to generate pure CW signals in SFT format, segmenting the data into time spans of $T_{\text{SFT}}=144{,}000$ s (4 hours) per segment. In practice, detector noise often exhibits non-Gaussian characteristics; however, for simplicity, we model it as Gaussian noise. This noise is generated directly in the Fourier domain, with the real and imaginary components of the Fourier transform defined by
@@
We use $\sqrt{S_{n}}/h_{0}$ to represent the noise level of the injected signal in our training and test data, as the ability to detect the signal depends on its amplitude relative to the detector’s noise amplitude. In this work, we consider only the first-order spindown term in Eq. (3). The parameter space defining the signals is listed in Table 1. We assess the neural network’s performance for an all-sky search with an observation time of T=921,600T=921{,}600 s ($\approx$ 10.67 days), chosen to approximate 10 days and to yield $2^{n}$ time segments to favor the architecture of the neural network.
@@ III Methods
Similar to (Dreissigacker and Prix, 2020; Joshi and Prix, 2024), we use the real and imaginary parts of SFT data as a two-dimensional input. Combining data from the H1 and L1 detectors, we obtain four channels: real and imaginary components for each detector. We perform a Fourier transform every 14,400 s on data spanning a total observation time of T=921,600T=921{,}600 s, yielding 64 time segments along the time axis. The frequency axis is cropped to retain 512 frequency bins, corresponding to a bandwidth of 35.56 mHz, sufficient to accommodate Doppler modulations due to Earth’s motion over the observation period. To enhance robustness, we then randomly shift the frequency axis, ensuring the CW signal remains within the selected range without always appearing at the center. The resulting preprocessed data have a shape of $(512,64,4)$, with dimensions selected as powers of 2 to favor the U-Net’s architecture. Finally, we normalize the data to the range [1,1][-1,1].
@@ III.4 Detection Statistic and Detectability
In addition to the loss function, we monitored the training process with $p_{\mathrm{det}}$, which measures the fraction of noisy CW signal samples correctly identified by the detection statistic at a fixed false alarm rate ($p_{\mathrm{fa}}$). Using Eq. 3 from (Joshi and Prix, 2024), we set $p_{\mathrm{fa}}$ to 1% per 50 mHz. For our data, spanning a 35.56 mHz frequency band per image, we calculated $p_{\mathrm{fa}}=0.71\%$. We selecte the model with the highest $p_{\mathrm{det}}$ on the validation set as the final model, since $p_{\mathrm{det}}$ is the metric for evaluating the model’s ability to identify CW signals in the presence of noise.
@@ IV Results
Figure 1: Loss (top) and detection probability $p_{\mathrm{det}}$ at a $1\%$ false alarm rate per $50\,\mathrm{mHz}$ (bottom) during the training process for the 20 Hz attention U-Net model on the training dataset (black) and validation dataset (colored) with different noise level $\sqrt{S_{n}}/h_{0}$.
@@ IV.1 Denoising
Figure 2: Denoising performance of the attention U-Net models trained at 20 Hz (left) and 500 Hz (right) on H1 spectrogram spanning 921,600 s. The attention U-Net takes a noisy input with noise level $\sqrt{S_{n}}/h_{0}=20\,\mathrm{Hz}^{-1}$ (top) and produces a denoised output (middle), which closely resembles the ground truth signal (bottom), recovering the major signal features.
@@
Figure 3: Denoising performance of the attention U-Net models trained at 20 Hz (left) and 500 Hz (right) on H1 spectrogram spanning 921,600 s with time gaps (duty factor = 0.75). Despite being trained on data without time gaps, the models successfully denoise the noisy input with noise level $\sqrt{S_{n}}/h_{0}=20\,\mathrm{Hz}^{-1}$ (top) to produce an output (middle) that closely resembles the ground truth signal (bottom), recovering the major signal features.
@@ IV.2 Sensitivity
We evaluated $p_{\mathrm{det}}$ across various noise levels $\sqrt{S_{n}}/h_{0}$ using 2,000 test samples (1,000 noisy signals and 1,000 pure noise) to determine the noise level at which a 90% detection probability is achieved ($\mathcal{D}^{90\%}$). Figure 4 shows the $p_{\mathrm{det}}$ of the 20 Hz model at different noise levels (blue curve). We also included the $p_{\mathrm{det}}$ of a matched-filter search using Weave from (Joshi and Prix, 2024), scaled to match the timespans in this study, as a benchmark (black curve). Our model achieves a sensitivity depth of $\mathcal{D}^{90\%}=29.97\pm 0.24\,\mathrm{Hz}^{-1/2}$. To compare with (Joshi and Prix, 2024), we scaled their $\mathcal{D}^{90\%}$ values using the relation $\mathcal{D}^{90\%}\propto T^{0.25}$, adjusting from their 10-day observation period to our 921,600 s ($\approx$ 10.67 days). This yields a sensitivity depth of $29.78\,\mathrm{Hz}^{-1/2}$ for their model. The sensitivity of our attention U-Net is comparable to the model from (Joshi and Prix, 2024) trained at 20 Hz.
Figure 4: Detection probability $p_{\mathrm{det}}$ (at a $1\%$ false alarm rate per $50\,\mathrm{mHz}$) versus noise level $\sqrt{S_{n}}/h_{0}$ at 20 Hz. The black curve shows the performance of the Weave matched-filter method, rescaled from (Joshi and Prix, 2024). The blue curve represents the attention U-Net’s performance, with a 90% sensitivity depth $\mathcal{D}^{90\%}=29.97\pm 0.24\,\mathrm{Hz}^{-1/2}$.
The performance of all models is summarized in Table 2, which reports the sensitivity depth $\mathcal{D}^{90\%}$ of our trained model, the rescaled sensitivity depth using the Weave matched-filter method $\mathcal{D}_{\mathrm{MF}}^{90\%}$ from (Joshi and Prix, 2024), and the detectability of our trained model at a noise level $\sqrt{S_{n}}/h_{0}=\mathcal{D}_{\mathrm{MF}}^{90\%}$. Our attention U-Net achieves comparable sensitivity at 20 Hz and 200 Hz but lower sensitivity at 500 Hz, 1000 Hz, and acroos the 20–1000 Hz range compared to (Joshi and Prix, 2024).
@@
Table 2: Sensitivity depth at 90% detection probability ($\mathcal{D}^{90\%}$) for the trained attention U-Net models, compared to the matched-filter sensitivity depth ($\mathcal{D}^{90\%}_{\mathrm{MF}}$) from (Joshi and Prix, 2024), and the detection probability of the trained attention U-Net at noise level $\sqrt{S_{n}}/h_{0}=\mathcal{D}_{\mathrm{MF}}^{90\%}$.
@@ IV.3 Generalization on Duty Factor and Total Observation Time
To study the dependence of sensitivity on the duty factor $\eta$, we evaluated the performance of the 20 Hz model on test data with time gaps, simulating realistic observation conditions. Figure 5 shows the variation of $\mathcal{D}^{90\%}$ with duty factors ranging from 0.5 to 1.0, where typical observation runs have duty factors of approximately 0.6–0.7. Fitting $\mathcal{D}^{90\%}$ to a power law reveals a scaling of $\mathcal{D}^{90\%}\propto\eta^{0.53\pm 0.02}$, with $\mathcal{D}^{90\%}=22.97\pm 0.26\,\mathrm{Hz}^{-1/2}$ at $\eta=0.6$.
Figure 5: Sensitivity depth $\mathcal{D}^{90\%}$ versus duty factor $\eta$ for the 20 Hz attention U-Net model trained on gap-free data. For $\eta=0.6$, the model achieves a sensitivity depth of $\mathcal{D}^{90\%}=22.97\pm 0.26\,\mathrm{Hz}^{-1/2}$, with a fitted power law indicating that $\mathcal{D}^{90\%}\propto\eta^{0.53\pm 0.02}$.
Our neural network was initially trained on data with $T\approx$ 10.67 days. To maximize sensitivity, we aim to use as much data as possible. Therefore, we investigate the impact of $T$ on the sensitivity by training the 20 Hz model with $T=$ 460,800 s, 921,600 s, 1,382,400 s, and 1,832,000 s (approximately 5, 10, 16, and 21 days, respectively). Figure 6 shows the resulting $\mathcal{D}^{90\%}$ values, achieving $\mathcal{D}^{90\%}=35.86\pm 0.38\,\mathrm{Hz}^{-1/2}$ at $T\approx$ 21 days. The sensitivity depth scales as $\mathcal{D}^{90\%}\propto T^{0.28\pm 0.01}$, similar to the $T^{0.25}$ scaling observed in semi-coherent search methods. Notably, the training time for our attention U-Net scales linearly with $T$, as the number of frequency bins is fixed. For sufficiently long $T$, such as a year of data, we have to increase the number of time bins by a factor of 35 and frequency bins by a factor of 7 to cover the entire signal for the parameter range defined in Table. 1. As a result, the overall scaling remains lower than that of semi-coherent searches, which scale as $\sim T^{2}$ for a fixed coherence time (Prix and Shaltev, 2012). This demonstrates that the attention U-Net architecture is feasible to scale up to longer observation time for better performance.
Figure 6: Sensitivity depth $\mathcal{D}^{90\%}$ versus total observation time $T$ for the 20 Hz attention U-Net model. For $T=1,832,000\,\mathrm{s}$ ($\approx$ 21 days), the model achieves a sensitivity depth of $\mathcal{D}^{90\%}=35.86\pm 0.38\,\mathrm{Hz}^{-1/2}$, with a fitted power law indicating that $\mathcal{D}^{90\%}\propto T^{0.28\pm 0.01}$.